Amy can run the 100 meter in 6 minutes. Betty can run the 100 meter in 8 minutes. If Betty goes out to practice and starts 1 minutes before Amy starts on her practice, will Amy catch up to Betty before she reaches the 100m marker?
Let x be the time in minutes for Amy to catch Betty.

Answer by jorel1380(3719)

100/8=12.5 m/min
100/6=16.67 m/min
12.5m+n(12.5)=n(16.67)
16.67n-12.5n=12.5
n=3 minutes before Amy catches Betty. Betty will have been running for 4 minutes by that time
☺☺☺☺